# The CCube reconstruction algorithm for the SoLid experiment

### v2

@@ 7 The SoLid approach
IBD events exhibit two characteristics that are useful for reconstruction. They are sparse, and the energy deposits from positron ionisation are larger than the two from the annihilation gamma ones. The former has to be reconstructed more accurately, since it is in the core of the antineutrino energy estimator. Therefore, starting with the assignment of energy to the most energetic cube before the others, an improvement in the performance of the reconstruction algorithm is expected; decreasing the ghost cube rate and lifting the degeneracy introduced in Figure 6.
The Orthogonal Matching Pursuit (OMP) algorithm fits this objective. At each stage, the aim is to find the best energy deposit that optimises the L2 norm (see Equation 4). The choice of the L2 norm over the KL norm is motivated by the simplicity of the implementation and the fact that it prioritises the most energetic cube in the event. Due to the higher energies and thus a larger number of photo avalanches, the Poisson behaviour is less prominent for these cubes. The method first adjusts the positron energy deposit before moving on to smaller energy deposits. Like FISTA, OMP will be used as the initializer of the ML-EM algorithm. In this case, the full capacity of the OMP is not required. To reduce computational cost, a simplified OMP (sOMP) is derived. The main difference between the two is that the simplified version allows one to assign the energy to a cube only once. Therefore, once a cube is accessed, it is removed from the list of options. Hence, the sOMP algorithm is run once and is used as an initialiser of ML-EM.
@@
The final comparison between the three methods is summarised in Table 4. The results presented are given for an energy threshold of 75 keV, which is an actual choice of the reconstuction adopted in SoLid.
@@ 7.1 Energy resolution
The combined sOMP and ML-EM approach shows dominant reconstruction performance. It is interesting to check that the energy resolution complies with the expectations of the detector design. This study is carried out using events selected by the so-called topological algorithm: empty 333 envelope around the most energetic cube (AC) in the event and two additional electromagnetic clusters reconstructed outside that envelope. Event candidates are also required to have the most energetic cube above 1.5 MeV. These conditions ensure that the AC candidate meets Equation1 requirements. Therefore, the energy spread is defined as:
where is the antineutrino energy from the Geant4 simulation and is the energy of the most energetic cube per ROSim event. Events in the sample defined above are split into 25 bins with identical statistics. For each bin, the distribution is built and fit with a Crystal Ball function crystalball . The fit result provides a value of , which is used to compute . In this case, is the median energy for each individual bin. The energy resolution studies for the same type of scintillator (EJ-200) are presented in ej200resolution . It has been chosen to adopt the same opportunistic 3-parameter function to fit the data:
@@ 8 Conclusion
The CCube reconstruction algorithm is presented in this article. It is of central importance for characterising the electromagnetic signal clusters in the SoLid reconstruction and therefore serves as the cornerstone of the definition and selection of IBD candidates recorded by the SoLid experiment. Three different methods employing established algorithms are designed to solve this reconstruction problem. The sOMP+EM-LM algorithm, which benefits from the use of the underlying physics of IBD events, exhibits the best performance. The choice of the 75 keV threshold was dictated by the original energy threshold in the antineutrino selection analysis. The combined sOMP + ML-EM algorithm is taken as the state-of-the-art method by the SoLid collaboration.

### v4

@@ 7 The SoLid approach
IBD events exhibit two characteristics that are useful for reconstruction. They are sparse, and the energy deposits from positron ionisation are larger than the two from the annihilation gamma ones. The former has to be reconstructed more accurately, since it is in the core of the antineutrino energy estimator. Therefore, starting with the assignment of energy to the most energetic cube before the others, an improvement in the performance of the reconstruction algorithm is expected. This improvement includes the decrease in the ghost cube rate and lift of the degeneracy introduced in Figure 6.
The Orthogonal Matching Pursuit (OMP) algorithm fits this objective. At each stage, the aim is to find the best energy deposit that optimises the L2 norm (see Equation 4). The choice of the L2 norm over the KL norm is motivated by the simplicity of the implementation and the fact that it prioritises the most energetic cube in the event. Due to the higher energies and thus a larger number of photo avalanches, the Poisson behaviour is less prominent for these cubes. The method first adjusts the positron energy deposit before moving on to smaller energy deposits. Like FISTA, OMP will be used as the initialiser of the ML-EM algorithm. In this case, the full capacity of the OMP is not required. To reduce computational cost, a simplified OMP (sOMP) is derived. The main difference between the two is that the simplified version allows one to assign the energy to a cube only once. Therefore, once a cube is accessed, it is removed from the list of options. Hence, the sOMP algorithm is run once and is used as an initialiser of ML-EM.
@@
The final comparison between the three methods is summarised in Table 4. The results presented are given for an energy threshold of 75 keV, which is the actual choice of the reconstruction adopted in SoLid.
@@ 7.1 Energy resolution
The combined sOMP and ML-EM approach grants the best reconstruction performance. It is interesting to check that the energy resolution complies with the expectations of the detector design. This study is carried out using events selected by the so-called topological algorithm: empty 333 envelope around the most energetic cube (AC) in the event and two additional electromagnetic clusters reconstructed outside that envelope. Event candidates are also required to have the most energetic cube above 1.5 MeV. These conditions ensure that the AC candidate meets the requirements of Equation 1. Therefore, the energy spread is defined as:
where is the antineutrino energy from the Geant4 simulation and is the energy of the cube with the highest energy in the ROSim event. Events in the sample defined above are split into 25 bins with identical statistics. For each bin, the distribution is built and fit with a Crystal Ball function [18]. The fit result provides a value of , which is used to compute . In this case, is the median energy for each individual bin. The energy resolution studies for the same type of scintillator (EJ-200) are presented in [19]. It has been chosen to adopt the same opportunistic 3-parameter function to fit the data:
@@
7.2 Open dataset performance
The reconstruction algorithm presented in this article and its performance have been developed and obtained using simulated data. The performance of the reconstruction algorithm on the data can be illustrated by a search for antineutrino signal in the SoLid detector. The reactor-off data sample and an open dataset are used (i.e. processed with the SoLid analysis software and the CCube reconstruction algorithm in particular) for this purpose. Reactor-on samples are recorded during periods when the BR2 reactor operates at nominal power. Hence, such samples contain both signal and background events. The open data set (denoted ROn) is a small reactor-on sample of 21 days recorded during a period of 19 June to 10 July 2018. Reactor-off samples are recorded during periods when the BR2 reactor is off, respectively. These samples are used to study background sources and model their properties. The reactor off data sample (ROff) used for this study is the sum of all reactor-off periods recorded during SoLid Phase I (mid-2018 to mid-2020). The IBD signal model is determined with simulation data. The sum of the signal and background models is then compared to the open data set.
The aim of the CCube algorithm is to reconstruct the position of a signal candidate cube and the energy deposit associated with it. The cube where positron annihilation occurs contributes the most to the antineutrino energy estimator. Therefore, the variables that evaluate the performance of the reconstruction are the energy of the annihilation cube itself (AC energy) and the spatial distance in cubes between the annihilation cube and the neutron capture cube (R, which serves as a measurement of the precision of position assignment).
The level of background in the raw SoLid data is overwhelming (the signal-to-background ratio is at the level of 1 to 1000). Antineutrino candidates are searched for in events where multiple cubes are present. This ensures that there is enough information to separate annihilation photons and positron signals. The CCube algorithm is instrumental in the separation of the annihilation gammas and positron ionisation components of the electromagnetic signal. This topological selection is followed by multivariate selections that take advantage of the physical characteristics of the signal candidates that can be built from cube information. For example, the angle between the momentum vectors of the annihilation gamma, built from the annihilation gamma and the positron cubes; a back-to-back radian angle will more likely designate a signal candidate.
Figure 9 shows the distributions of the energy of the positron candidates and their distance from the neutron-capture cube candidate in the ROn sample. The sum of the histograms of the background (red) and signal (grey) models agrees with the open data set (blue data points) within statistical uncertainty. There is a clear excess of events with respect to the ROff prediction, consistent with antineutrino IBD candidates. The reconstruction algorithm described in this article is primarily used to define the signal. It also proved to be essential in the definition of discriminative variables in order to successfully select the signal.
@@ 8 Conclusion
The CCube reconstruction algorithm is presented in this article. It is of central importance for characterising the electromagnetic signal clusters in the SoLid reconstruction and therefore serves as the cornerstone of the definition and selection of IBD candidates recorded by the SoLid experiment. Three different methods employing established algorithms are designed to solve this reconstruction problem. The sOMP+ML-EM algorithm, which benefits from the use of the underlying physics of IBD events, exhibits the best performance. The choice of the 75 keV threshold was dictated by the original energy threshold in the antineutrino selection analysis. The combined sOMP+ML-EM algorithm is considered the state-of-the-art method by the SoLid collaboration.